OpenStudy (anonymous):

Please help me with the question : show that Q[x]/(x^2-2) is not isomorphic with Q[x]/(x^2-3)

OpenStudy (misty1212):

hmm i think i remember this gimmick but it was with \(\mathbb{Q}(\sqrt2)\) and \(\mathbb{Q}(\sqrt3)\)

OpenStudy (misty1212):

since they are the same thing, it must be the same the hint is to show that if you had such an isomorphims, since \(\phi(1)=1\) it must be the case that \(\phi(\sqrt2)=\sqrt2\) but \(\sqrt 2\notin \mathbb{Q}(\sqrt3)\)

OpenStudy (anonymous):

you are right it is similar , but this one is little bit more trickier!